# On the Hardness of Approximating Distributions with Probabilistic Circuits

### v1

@@ 3 Approximate Modeling with Tractable Marginals is $\mathsf{NP}$-hard
Given a probability distribution $P$ and a $k$-convex $f$-divergence $D_{f}$, for any $0\!<\!\epsilon\!<\!\frac{1}{4}$, it is $\mathsf{NP}$-hard to represent its $k\epsilon^{2}$-$D_{f}$-approximation as a model that admits tractable marginals.
For the full proof, see Appendix E.2. We prove this using a reduction to SAT. First, we construct a formula $f$ over $\mathbf{X}$, and another formula $f^{\prime}$ over $\mathbf{X}$ and $Y$ where $f^{\prime}=(Y\wedge f)\vee(\neg Y\wedge X_{1}\wedge\cdots\wedge X_{n}).$ Let $P$ be uniform over the models of $f^{\prime}$ and $Q$ be a model such that $D_{f}(P\|Q)<k\epsilon^{2}$ with $0\leq\epsilon<\frac{1}{4}$. Then, using bounded difference between marginals, we show that $f$ is satisfiable if and only if $Q(Y=1)\geq\frac{1}{4}.$
@@
Given a probability distribution $P$, for $0<\epsilon<\frac{1}{4}$, it is $\mathsf{NP}$-hard to represent its $\epsilon$-$D_{\mathsf{TV}}$-approximation as a model that can tractably compute marginals.
@@
Thus, the class of polynomial-sized smooth and decomposable PCs, cannot contain $k\epsilon^{2}$-$\mathcal{D}_{f}$-approximations for all distributions unless $\mathsf{P}=\mathsf{NP}$. While Martens and Medabalimi [26] previously showed a related result that there exists a function for which a sequence of decomposable PCs converging to approximate the function arbitrarily well requires an exponential size, our result applies more broadly to any class of models supporting tractable marginals as well as allowing for bounded but non-vanishing approximation error. Thus, representing functions or distributions as decomposable PCs, even approximately, remains hard, challenging the feasibility of approximate compilation for approximate inference with small error.
4 Large Dec. & Det. PCs for Approximate Modeling
@@
Continuing our characterization of approximation power of PCs, we now turn to the family of decomposable and deterministic PCs. We take inspiration from related results for logical circuits. In particular, Bova et al. [1] proved an exponential separation of DNNFs from d-DNNFs; [^1] Decomposable negation normal forms (DNNFs) and deterministic, decomposable negation normal forms (d-DNNFs). De Colnet and Mengel [18] showed that there exists functions which require exponentially sized d-DNNF to approximate.
@@
Nevertheless, this does not immediately imply the same separation for probabilistic circuits for two main reasons: (1) approximation for PCs is measured in terms of divergences between distributions rather than some probabilistic error between Boolean functions, and (2) our approximator can represent arbitrary distributions instead of being limited to a Boolean function (or a uniform distribution over it). This section presents our proof of exponential separation of decomposable PCs from decomposable & deterministic ones, by constructing a family of distributions that can be represented by compact decomposable PCs but any PC that is also deterministic and approximates it within a bounded TV distance requires exponential size.
@@
We consider the Sauerhoff function [34] which was used to show the separation between DNNFs and d-DNNFs for exact compilation [1]. Let $g_{n}:\{0,1\}^{n}\to\{0,1\}$ be a function evaluating to 1 if and only if the sum of its inputs is divisible by 3. The Sauerhoff function is defined as $S_{n}:\{0,1\}^{n^{2}}\to\{0,1\}$ over the $n\times n$ matrix $X=(x_{ij})_{1\leq i,j\leq n}$ by $S_{n}(X)=R_{n}(X)\vee C_{n}(X)$ where $R_{n},C_{n}:\{0,1\}^{n^{2}}\to\{0,1\}$ are defined by $R_{n}(X)=\bigoplus_{i=1}^{n}g_{n}(x_{i,1},x_{i,2},\ldots,x_{i,n})$ and $C_{n}(X)=R_{n}(X^{T})$. Here, $\oplus$ represents addition modulo 2.
@@
There exists a DNNF of size $O(n^{2})$ that exactly represents the Sauerhoff function $S_{n}$, constructed using two compact ordered binary decision diagrams to represent $R_{n},C_{n}$ [1, Proposition 7]. We then define our family of target distributions $P_{n}$ as follows: let $\mathcal{C}_{n}$ be a DNNF for $S_{n}$ with size $O(n^{2})$; then $P_{n}$ is a decomposable PC obtained by replacing the literals of $\mathcal{C}_{n}$ with corresponding indicator functions, $\vee$ with $\oplus$, and $\wedge$ with $\otimes$ nodes with uniform parameters, then smoothing. Note that $P_{n}(\mathbf{x})>0$ if and only if $S_{n}(\mathbf{x})=1$. We will show that a deterministic, decomposable PC approximating $P_{n}$ requires exponential size.
@@
A deterministic, decomposable PC that is a $\epsilon$-$D_{\mathsf{TV}}$-Approximator of $P_{n}$, where $\epsilon=\frac{1}{16}-\Omega(1/\text{Poly}(n^{2}))$, has size $2^{\Omega(n)}$.
@@
We say that a Boolean formula $g$ is a weak $\epsilon$-approximation of another Boolean formula $f$ if $\operatorname{MC}(f\wedge\neg g)+\operatorname{MC}(\neg f\wedge g)\leq\epsilon 2^{n}.$
@@
Let $0\leq\epsilon<\frac{1}{8}$ and $P$ be a uniform distribution whose support is given by a Boolean function $f$. Suppose that $Q$ is a decomposable and deterministic PC representing an $\epsilon$-$D_{\mathsf{TV}}$-Approximator of $P$. Then there exists a d-DNNF $g$ which has size polynomial in the size of $Q$ that represents a $4\epsilon$-weak-approximator of $f$.
For full proof see Appendix E.4. Start with our circuit $Q$ with support $g$, and use Algorithm E.3 to prune $Q(\mathbf{x})<\frac{1}{2^{n+1}}$. We call this unnormalized distribution $Q^{\prime}$ and it’s support $g^{\prime}$. Using $D_{\mathsf{TV}}(P\|Q)<\epsilon$, the fact that on $g^{\prime}$ all $Q(\mathbf{x})\geq\frac{1}{2^{n+1}}$, and that $\operatorname{MC}(f)\leq 2^{n}$, we show that we can then bound $\operatorname{MC}(f\wedge\neg g^{\prime})+\operatorname{MC}(\neg f\wedge g^{\prime})<4\epsilon 2^{n}$, which corresponds to a $4\epsilon$-weak-approximation.
@@
Let $\mathcal{C}$ be a d-DNNF that is a $(\frac{1}{4}-\Omega(1/\text{Poly}(n^{2})))$-weak-approximation of $S_{n}$. Sauerhoff [34] showed that any “two-sided” rectangle approximation [^2] See Appendix D for details on rectangle partitions (matching the notion of weak approximation) of $S_{n}$ within $\frac{1}{4}-\Omega(1/\text{Poly}(n^{2}))$ must have size $2^{\Omega(n)}$. Bova et al. [1] further showed that a d-DNNF $\mathcal{C}$ computing a function $f$ is a balanced rectangle partition of $f$ with size at most $\left\lvert\mathcal{C}\right\rvert$. Thus, $\left\lvert\mathcal{C}\right\rvert=2^{\Omega(n)}$. ∎
Since by we can represent $S_{n}$ by a DNNF of size $O(n^{2})$, there exists an exponential gap in approximation between DNNF and d-DNNF. We next show this exponential gap in approximation persists for PCs.
  • Proof Sketch of Theorem 4.1.
@@
    See Appendix E.5 for the full proof. Suppose we have a deterministic, decomposable PC $Q$ that is an $\epsilon$-$D_{\mathsf{TV}}$-Approximator of $P_{n}$ for $\epsilon\!=\!(\frac{1}{16}\!-\!\Omega(1/Poly(n^{2})))$. Consider the uniform distribution $U$ over $S_{n}$. We use the triangle inequality to bound $D_{\mathsf{TV}}(U\|Q)\leq D_{\mathsf{TV}}(U\|P_{n})\!+\!D_{\mathsf{TV}}(P_{n}\|Q)$. Given that $D_{\mathsf{TV}}(P_{n}\|Q)\!<\!\frac{1}{16}\!-\!\Omega(1/Poly(n^{2})))$, it is only left to bound $D_{\mathsf{TV}}(U\|P_{n})$. Taking the d-DNNF constructed by [1, Prop. 7] and the low 0-density property of the Sauerhoff function [34], we bound $D_{\mathsf{TV}}(P_{n}\|U)<\eta$ for $\eta\!=\!\frac{1}{(1-1/\sqrt{2})2^{n^{2}}}$, which implies $D_{\mathsf{TV}}(U\|Q)<\epsilon-\Omega(1/Poly(n^{2}))+\eta$. We combine $\Omega(1/Poly(n^{2}))+\eta$ to simplify this expression. By Proposition 4.3, we can construct a d-DNNF $\mathcal{C}^{\prime}$ from $Q$—replacing $\oplus$ and $\otimes$ with $\vee$ and $\wedge$ respectively—that is a ($\frac{1}{4}\!-\!\Omega(1/Poly(n^{2}))$)-weak-approximation of $S_{n}$. By Proposition 4.4, $\left\lvert Q\right\rvert\!=\!\left\lvert\mathcal{C}^{\prime}\right\rvert\!=\!2^{\Omega(n)}$. ∎
@@
As we have constructed a decomposable PC $P_{n}$ that has size $O(n^{2})$ such that any deterministic, decomposable PC approximating it has size $O(2^{\Omega(n)})$, we have shown an unconditional exponential gap for approximation between decomposable PCs and deterministic, decomposable PCs. Thus, approximate modeling does not grant the ability to overcome exponential expressive efficiency gaps.
5 Conclusion
We established the hardness of approximating distributions with probabilistic circuits such that the $f$-divergence is small. First, we showed that this task is $\mathsf{NP}$-hard for any model that is tractable for marginal inference, including decomposable PCs. Then, we used the Sauerhoff function to demonstrate an exponential gap between decomposable PCs and deterministic, decomposable PCs when allowing for bounded approximation error. This proves that the succinctness gap in the case of exact compilation persists even under relaxed approximation conditions.
These results highlight key challenges in learning compact and expressive PCs while maintaining tractable inference. In light of this, we ask: can a polynomial-time algorithm enable learning an $\epsilon$-approximator for a broad family of distributions with a more relaxed $\epsilon$? Furthermore, does there exist modeling conditions that are sufficient for relative approximation of various queries? Lastly, we see this work as a first step to encourage further theoretical studies on approximate modeling and inference with guarantees using tractable models. In this paper, we focused on PCs that are tractable for marginal and MAP queries, but there are large classes of tractable models whose efficient approximation power remain unknown.
Appendix A Background
A.1 Probabilistic Circuits
Definition A.1 (Probabilistic circuits).
A probabilistic circuit (PC) $\mathcal{C}:=(\mathcal{G},\theta)$ represents a joint probability distribution $p(\mathbf{X})$ over random variables $\mathbf{X}$ through a directed acyclic graph (DAG) $\mathcal{G}$ parameterized by $\theta$. The DAG is composed of 3 types of nodes: leaf, product $\otimes$, and sum nodes $\oplus$. Every leaf node in $\mathcal{G}$ is an input, Boolean indicator functions in this paper, and every internal node receives inputs from its children $\operatorname{in}(n)$. The scope of a given node, $\phi(n)$, is the set of random variables as inputs to $n$. Each node $n$ of a PC is recursively defined as:
$p_{n}(\mathbf{x}):=\begin{cases}l(x),&\text{ if $n$ is a leaf}\\ \prod_{c\in\operatorname{in}(n)}p_{c}(\mathbf{x})&\text{ if $n$ is a $\otimes$}\\ \sum_{c\in\operatorname{in}(n)}\theta_{n,c}p_{c}(\mathbf{x})&\text{ if $n$ is a $\oplus$}\\ \end{cases}$ (1)
where $\theta_{n,c}\in[0,1]$ is the parameter associated with the edge connecting nodes $n,c$ in $\mathcal{G}$, and $\sum_{c\in\operatorname{in}(n)}\theta_{n,c}=1.$ The distribution represented by the circuit is the output at its root node.
[…]
A key characteristic of PCs is that imposing certain structural properties on the circuit enables tractable (polytime) computation of various queries. In this paper we focus on two families of PCs: those that are tractable for marginal inference and for maximum-a-posteriori (MAP) inference. The class of marginal queries of a joint distribution $p(\mathbf{X})$ over variables $\mathbf{X}$ refers the set of functions that can compute $p(\mathbf{y})$ for some assignment $\mathbf{y}$ for $\mathbf{Y}\subseteq\mathbf{X}$. Marginalization is a fundamental statistical operation which enables reasoning about subsets of variables, essential for tasks like decision making, learning, and predicting under uncertainty. While marginal inference is #$\mathsf{\#P}$-hard in general [32, 11], the family of PCs satisfying certain structural conditions, admit tractable marginal inference—specifically in linear time in the size of the circuit [15].
Definition A.2 (Smoothness and Decomposability).
A sum unit is smooth if its children have identical scopes: $\phi(c)=\phi(n),\forall c\in\operatorname{in}(n)$. A product unit is decomposable if its children have disjoint scopes: $\phi(c_{i})\cap\phi(c_{j})=\emptyset,\forall c_{i},c_{j}\in\operatorname{in}(n),i\neq j$. A PC is smooth and decomposable iff every sum unit is smooth and every product unit is decomposable.
Definition A.3 (Determinism).
A sum node is deterministic if, for any fully-instantiated input, the output of at most one of its children is nonzero. A PC is deterministic iff all of its sum nodes are deterministic.
Appendix B Relationship between Approximate Modeling and Inference

### v2

@@ 3.2 Hardness of Approximating Distributions using Tractable Models for Marginals
Given a (potentially unnormalized) probability distribution $\hat{P}$ and a $k$-convex $f$-divergence $D_{f}$, for any $0<\epsilon<\frac{1}{4}$, it is $\mathsf{NP}$-hard to represent the $k\epsilon^{2}$-$D_{f}$-approximation of its normalized distribution $P$ as a model that can tractably compute marginals.
Proof.
We will prove the above using a reduction from SAT. Let $\hat{P}$ be a Boolean formula over $\mathbf{X}=\{X_{1},\ldots,X_{n}\}$ and $\epsilon<\frac{1}{4}$. See that $\hat{P}$ satisfies all requirements to be considered an unnormalized probability distribution. We then define a new Boolean formula $\hat{P}^{\prime}$ over $\mathbf{X}$ and an auxiliary variable $Y$: $\hat{P}^{\prime}=(Y\wedge\hat{P})\vee(\neg Y\wedge X_{1}\wedge\cdots\wedge X_{n})$. Clearly $\hat{P}^{\prime}$ has $\operatorname{MC}(\hat{P})+1$ models. Let us now define a uniform distribution $P$ over these models of $\hat{P}^{\prime}$ (i.e., by normalizing $\hat{P}^{\prime}$). Suppose that we can efficiently obtain a probability distribution $Q$ such that $D_{f}(P\|Q)<k\epsilon^{2}$, which in turn implies that $D_{\mathsf{TV}}(P\|Q)<\epsilon$. From the definition of total variation distance, $\left\lvert P(Y=1)-Q(Y=1)\right\rvert<\epsilon<\frac{1}{4}$.
[…]
$Q(Y=1)$ $>\frac{\operatorname{MC}(\hat{P})}{1+\operatorname{MC}(\hat{P})}-\frac{1}{4}\geq\frac{1}{2}-\frac{1}{4}\geq\frac{1}{4}.$
Therefore, $\hat{P}$ is satisfiable if and only if $Q(Y=1)\geq\frac{1}{4}$. In other words, we can decide SAT if we can efficiently compute an $k\epsilon^{2}$-$D_{f}$-approximation as a model that supports tractable marginals. ∎
The following corollary immediately follows from the above proof.
@@
Given a (potentially unnormalized) probability distribution $\hat{P}$, for $0<\epsilon<\frac{1}{4}$, it is $\mathsf{NP}$-hard to represent the $\epsilon$-$D_{\mathsf{TV}}$-approximation of its normalized distribution $P$ as a model that can tractably compute marginals.
@@
Thus, the class of polynomial-sized probabilistic models supporting tractable marginals cannot contain $k\epsilon^{2}$-$\mathcal{D}_{f}$-approximations for all distributions unless $\mathsf{P}=\mathsf{NP}$. For instance, this includes decomposable PCs [8], sum of squares circuits [31], probabilistic generating circuits [50], determinantal point processes [1], Inception PCs [47], and positive unital circuits [22]. Furthermore, by Pinsker’s inequality this suggests that there exists distributions for which obtaining a decomposable PC with bounded KL-divergence of $1/8$ is $\mathsf{NP}$-hard. While Martens and Medabalimi [33] previously showed a related result that there exists a function for which a sequence of decomposable PCs converging to approximate the function arbitrarily well requires an exponential size, our result applies more broadly to any class of models supporting tractable marginals as well as allowing for a bounded but non-vanishing approximation error. Thus, it is difficult not only to exactly represent functions or distributions as compact decomposable PCs, but also to approximate within some small distance. Consequently, this presents a major challenge to the idea of using approximate compilation of PCs for approximate inference if the desired error tolerance $\epsilon$ is sufficiently small.
4 Large Deterministic & Decomposable PCs for Approximate Modeling
@@
Continuing our characterization of the approximation power of PCs, we now turn to the family of deterministic and decomposable PCs. We take inspiration from related results for logical circuits. In particular, Bova et al. [2] proved an exponential separation between DNNFs and d-DNNFs: i.e., there is a family of Boolean functions that can be compactly represented by decomposable circuits but requires exponentially sized deterministic and decomposable circuits. Furthermore, De Colnet and Mengel [21] showed that there exist functions that require exponential size to approximate with d-DNNFs under two notions of approximation for Boolean functions.
@@
Nevertheless, this does not immediately imply the same separation for probabilistic circuits due to two key reasons: (1) approximation for PCs is measured in terms of divergences between distributions rather than some probabilistic error between Boolean functions, and (2) our approximator can represent arbitrary distributions instead of being limited to a Boolean function (or a uniform distribution over it). This section presents our proof of exponential separation between the class of decomposable PCs and that of deterministic and decomposable ones, by constructing a family of distributions that can be represented by compact decomposable PCs, but any PC that is also deterministic and approximates it within a bounded TV distance must have an exponential size.
@@
We consider the Sauerhoff function [41] which was used to show the separation between DNNFs and d-DNNFs for exact compilation [2]. Let $g_{n}:\{0,1\}^{n}\to\{0,1\}$ be a function evaluating to 1 if and only if the sum of its inputs is divisible by 3. The Sauerhoff function is defined as $S_{n}:\{0,1\}^{n^{2}}\to\{0,1\}$ over the $n\times n$ matrix $X=(x_{ij})_{1\leq i,j\leq n}$ such that $S_{n}(X)=R_{n}(X)\vee C_{n}(X),$ where $R_{n},C_{n}:\{0,1\}^{n^{2}}\to\{0,1\}$ are defined as $R_{n}(X)=\bigoplus_{i=1}^{n}g_{n}(x_{i,1},x_{i,2},\ldots,x_{i,n})$ and $C_{n}(X)=R_{n}(X^{T})$. Here, $\oplus$ represents addition modulo 2.
@@
There exists a DNNF of size $O(n^{2})$ that exactly represents the Sauerhoff function $S_{n}$, constructed as a disjunction of two compact ordered binary decision diagrams (OBDDs)—a more restrictive kind of deterministic and decomposable circuits—that represent $R_{n}$ and $C_{n}$, respectively [2, Proposition 7]. We then define our family of target distributions $P_{n}$ as follows: let $\mathcal{C}_{n}$ be a DNNF for $S_{n}$ with size $O(n^{2})$; then $P_{n}$ is a decomposable PC obtained by replacing the literals of $\mathcal{C}_{n}$ with corresponding indicator functions, $\wedge$ with $\otimes$, and $\vee$ with $\oplus$ nodes with uniform parameters, followed by smoothing the circuit. [^6] Smoothing a decomposable PC takes polynomial (worst-case quadratic) time [8]. Note that $P_{n}$ outputs a positive value on an input $\mathbf{x}$ if and only if $S_{n}(\mathbf{x})=1$. We will show that a deterministic and decomposable PC approximating $P_{n}$ requires exponential size.
@@
A deterministic and decomposable PC that is a $\epsilon$-$D_{\mathsf{TV}}$-approximator of $P_{n}$ for some $\epsilon\leq\frac{1}{16}-\Omega(1/\text{Poly}(n^{2}))$ has size $2^{\Omega(n)}$.
@@
A Boolean formula $g$ is a weak $\epsilon$-approximation of another Boolean formula $f$ if $\operatorname{MC}(f\wedge\neg g)+\operatorname{MC}(\neg f\wedge g)\leq\epsilon\cdot 2^{n}.$
@@
Let $0\leq\epsilon<\frac{1}{8}$ and $P$ be a uniform distribution whose support is given by a Boolean function $f$. Suppose that $Q$ is a deterministic and decomposable PC representing an $\epsilon$-$D_{\mathsf{TV}}$-approximator of $P$. Then there exists a d-DNNF $g$ which has size polynomial in the size of $Q$ that represents a $4\epsilon$-weak-approximator of $f$.
Proof.
Suppose that $P$ is a uniform distribution over the support given by a Boolean function $f$, and $Q$ a probability distribution such that $D_{\mathsf{TV}}(P\|Q)<\epsilon<\frac{1}{8}$. Then, we can construct an (unnormalized) deterministic and decomposable PC $Q^{\prime}$ by pruning the edges of $Q$ such that any assignment $\mathbf{x}$ is eliminated from the resulting support of $Q^{\prime}$ if and only if $Q(\mathbf{x})<\frac{1}{2^{n+1}}$. We call this support $g$. We provide this pruning algorithm, which relies on determinism as a key property, in Appendix A.3.2. Clearly, the size of $Q^{\prime}$ is at most the size of $Q$. We will now briefly summarize how this pruning scheme induces a weak approximation and refer to Appendix A.3.3 for the full derivation. Intuitively, we know that the support of $Q$ must cover most of the support of $P$, as $P$ is a uniform distribution and $D_{\mathsf{TV}}(P\|Q)$ is bounded. While it is possible for the support of $Q$ to be much larger than the support of $P$, then the probability assigned by $Q$ to assignments outside of the support of $P$ must also be very small to maintain a small TV distance. Thus, pruning away these assignments with small probability allows us to retrieve $Q^{\prime}$ whose support has only a small number of non-overlapping assignments with models of $f$. More formally, we know that by our pruning scheme, $Q(\mathbf{x})\geq\frac{1}{2^{n+1}}$ on $g$ and $Q(\mathbf{x})<\frac{1}{2^{n+1}}$ on $\neg g$. Using this fact, we can lower bound our original total variation distance by $\frac{1}{2}(\operatorname{MC}(f\wedge\neg g)/2^{n+1}+\operatorname{MC}(\neg f\wedge g)/2^{n+1}$), which implies that $\operatorname{MC}(f\wedge\neg g)+\operatorname{MC}(\neg f\wedge g)<4\epsilon\cdot 2^{n}$. Thus, $g$ is a $4\epsilon$-weak-approximator of $f$, and we can represent it as a polynomially sized d-DNNF by taking the deterministic and decomposable PC $Q^{\prime}$ and converting it to a logical circuit. ∎
@@
Let $\mathcal{C}$ be a d-DNNF such that it is a $(\frac{1}{4}-\Omega(1/\text{Poly}(n^{2})))$-weak-approximation of $S_{n}$. Sauerhoff [41] showed that any “two-sided” rectangle approximation [^7] See Appendix A.2 for details on rectangle partitions. (which matches the notion of weak approximation) of $S_{n}$ within $\frac{1}{4}-\Omega(1/\text{Poly}(n^{2}))$ must have size $2^{\Omega(n)}$. Bova et al. [2] further showed that a d-DNNF $\mathcal{C}$ computing a function $f$ is a balanced rectangle partition of $f$ with size at most $\left\lvert\mathcal{C}\right\rvert$. Thus, $\mathcal{C}$ must have size $2^{\Omega(n)}$. ∎
We are now ready to prove our main result about exponential size lower bound on deterministic and decomposable PCs as approximators.
Proof of Theorem 4.1.
@@
Suppose that we have a deterministic and decomposable PC $Q$ that is an $\epsilon$-$D_{\mathsf{TV}}$-approximator of $P_{n}$, where $\epsilon=(\frac{1}{16}-\Omega(1/Poly(n^{2})))$. Consider then the uniform distribution $U$ over $S_{n}$. Then, by the triangle inequality, $D_{\mathsf{TV}}(U\|Q)\leq D_{\mathsf{TV}}(U\|P_{n})+D_{\mathsf{TV}}(P_{n}\|Q)<D_{\mathsf{TV}}(U\|P_{n})+\epsilon.$ By Bova et al. [2, Proposition 7], we know that the DNNF constructed to represent the Sauerhoff function is a disjunction of two OBDDs, which respectively represent $R_{n},C_{n}$. As each OBDD can easily be translated to a d-DNNF with a polynomial size increase, their PC counterparts (OR to $\oplus$ and AND to $\otimes$) will still represent the same Boolean functions [6]. Thus, the non-deterministic PC representing $P_{n}$ based on this construction only has one non-deterministic sum node at the root, and can return values at most 2. This allows us to see that $D_{\mathsf{TV}}(U\|P)\leq\left\lvert\frac{1}{\left\lvert S_{n}\right\rvert}-\frac{2}{\left\lvert S_{n}\right\rvert+1}\right\rvert<\frac{1}{\left\lvert S_{n}\right\rvert}<\eta$ where $\eta=\frac{1}{(1-1/\sqrt{2})2^{n^{2}}}$ based on the fact that $\left\lvert S_{n}\right\rvert>(1-\beta)2^{n^{2}}$ for $\beta<1/\sqrt{2}$, derived from the low 0-density property of $S_{n}$ under the uniform distribution [41]. Therefore, $D_{\mathsf{TV}}(U\|Q)<\epsilon+\eta$. Since, the $\Omega(1/Poly(n^{2}))$ term in $\epsilon$ subsumes $\eta$, we can more simply say $D_{\mathsf{TV}}(U\|Q)<\frac{1}{16}-\Omega(1/Poly(n^{2}))$. Next, we construct a d-DNNF $\mathcal{C}^{\prime}$ from $Q$ by replacing indicators with literals, $\oplus$ with $\vee$, and $\otimes$ with $\wedge$. By Proposition 4.3, $\mathcal{C}^{\prime}$ is a $(\frac{1}{4}-\Omega(1/Poly(n^{2})))$-weak-approximation of $S_{n}$. Thus, by Proposition 4.4, $\left\lvert Q\right\rvert=\left\lvert\mathcal{C}^{\prime}\right\rvert=2^{\Omega(n)}$. ∎
@@
To sum up, we constructed a decomposable PC $P_{n}$ that has size $O(n^{2})$ such that any deterministic and decomposable PC approximating it has size $O(2^{\Omega(n)})$, thereby showing an unconditional exponential gap for approximation between decomposable PCs and deterministic and decomposable PCs. This result highlights a fundamental limitation: approximate modeling does not grant us an additional flexibility to overcome exponential expressive efficiency gaps. Moreover, we next show that approximate modeling, even when somehow obtained, is unfortunately still not enough to use PCs for efficient approximate inference with guarantees in general.
5 Relationship between Approximate Modeling and Inference